# Skill-aware Mutual Information Optimisation for Generalisation in Reinforcement Learning

### v1

@@ D.3 Implementation details
In this section, we provide the implementation details for SaMI. Our codebase is built on top of the publicly released implementation Stable Baselines3 by Raffin et al. [2021] as well as the implementation of InfoNCE by Oord et al. [2019]. A public and open-source implementation of SaMI is available at https://github.com/uoe-agents/SaMI.
Base algorithm. We use SAC [Haarnoja et al., 2018] for downstream evaluation of the learned context embedding. SAC is an off-policy actor-critic method that uses the maximum entropy framework for soft policy iteration. At each iteration, SAC performs soft policy evaluation and improvement steps. We use the same SAC implementation across all baselines and other methods. We trained SAC agents for 1.6 million timesteps on Panda-gym and MuJoCo benchmarks (i.e. Half-cheetah, Ant, Crippled Half-cheetah, Crippled Ant, SlimHumanoid). We evaluated trained agents every $10,000$ timesteps on tasks with fixed random seeds.
Table 4: Hyperparameters used in Panda-gym and MuJoCo benchmarks. Most hyperparameter values are unchanged across tasks except for contrastive batch size and SaNCE loss coefficient.
@@
Encoder architecture. For our method, the context encoder $\psi$ is modelled as a Long Short-Term Memory (LSTM) that produces a 128-dimensional hidden state vector, subsequently processed through a single-layer feed-forward network to generate a 6-dimensional context embedding. The actor and critic both use the same context encoder to embed trajectories. For contrastive learning, SaNCE utilises a momentum encoder $\psi^{*}$ to generate positive and negative context embeddings [Laskin et al., 2020, He et al., 2020]. Formally, denoting the parameters of $\psi$ as $\theta_{\psi}$ and those of $\psi^{*}$ as $\theta_{\psi^{*}}$, we update $\theta_{\psi^{*}}$ by:
@@
Here $m\in[0,1)$ is a soft-update rate. Only the parameters $\theta_{\psi}$ are updated by back-propagation. The momentum update in Equation 14 makes $\theta_{\psi^{*}}$ evolve more smoothly by having them slowly track the $\theta_{\psi}$ with $m\ll 1$ (e.g., $m=0.05$ in this research). This means that the target values are constrained to change slowly, greatly improving the stability of learning.
@@ E.1 Balance contrastive and RL updates: loss coefficient $\alpha$
While past work has learned hyperparameters to balance the contrastive loss coefficient $\alpha$ relative to the RL objective [Jaderberg et al., 2016, Bachman et al., 2019], we use both the contrastive and RL objectives together with equal weight $\alpha=1.0$ to be optimal for the MuJoCo benchmark, and $\alpha=0.01$ for the Panda-gym benchmark, we also analyse the effect of loss coefficient $\alpha$ for CCM, SaTESAC and SaCCM in MuJoCo (Figure 11) and Panda-gym (Figure 10) benchmarks.
@@ E.2.1 Buffer size
Figure 12: Comparison of different buffer sizes in MuJoCo benchmark in training tasks (over 5 seeds). Buffer size = 400000, 100000, 10000, and 1000.
@@
Figure 13: Comparison of different contrastive batch sizes in MuJoCo benchmark in training tasks (over 5 seeds). contrastive batch size = 512, 128, 16, and 8.
@@ F.1.1 Visualisation of context embedding
We visualise the context embedding via t-SNE [Van der Maaten and Hinton, 2008] (Figure 15) and PCA [Jolliffe, 2002] (Figure 14). We found that when the mass of the cube is high (30 Kg and 10 Kg), the agent learned the Push skill (the yellow bounding box in Figure 5.1(a)), whereas with lower masses, the agent learned the Pick&Place skill. However, as shown in Figure 15(b), CCM did not display clear skill grouping. This indicates that SaMI extracts high-quality skill-related information from the trajectories and helps with embodying diverse skills.
Figure 14: PCA visualization of context embedding extracted from trajectories collected in Panda-gym environments.
Figure 15: t-SNE visualization of context embedding extracted from trajectories collected in Panda-gym environments.
@@ F.1.2 Heatmap of Panda-gym benchmark
Figure 16: Heatmap of success rate and learned skills of SaCCM. For each grid, the $(i,j)^{th}$ location shows the probability of the skills executed in 100 evaluations with $(\text{mass}=i,\text{friction}=j)$. We determine the skills by detecting contact points between the end effector and the cube and between the cube and the table.
This section provides the heatmap results and further analysis of TESAC, CCM, and SaTESAC on the Panda-gym benchmark. Initially, from the heatmap results of SaTESAC and SaCCM (Figure 6), we observed that with higher cube masses (30 and 10 kg), the agent adopted the Push skill (as indicated by the clustered points in the yellow bounding box in Figure 5.1(a)). At lower masses, the agent adopted the Pick&Place skill.
In contrast to CCM, as seen in Figures 16(e-f), CCM predominantly learned the Pick&Place skill, resulting in a decline in success rates for tasks at $mass=30$, as the agent could not lift the cube off the table using the Pick&Place skill, as depicted in Figure 16(d). The visualisation of the context embedding (Figure 15) did not show clear grouping across different tasks.
Finally, TESAC only mastered the Push skill. The Push skill is relatively simpler to learn than the Pick&Place skill, as it does not require the agent to master manipulating fingers to pick up cubes. Consequently, TESAC’s success rate notably decreased in environments with higher friction.
@@ F.2 MuJoCo
We find that SaMI help the RL agents be versatile and embody multiple skills. For example, Figure 17 displays the t-SNE visualization results of four methods in the Crippled Half-cheetah environment for both test (moderate and extreme) and training tasks. Tasks corresponding to different crippled joints on the same leg are clustered together. Further, we have displayed rendered videos (https://github.com/uoe-agents/SaMI) to better demonstrate the different skills learned.
Figure 17: t-SNE and PCA visualisation of context embedding extracted from trajectories collected in Crippled Half-cheetah environment.
Figure 18: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Crippled Ant environment.
Figure 19: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Half-cheetah environment.
Figure 20: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Ant environment.
Figure 21: t-SNE and PCA visualization of context embedding extracted from trajectories collected in SlimHumanoid environment.
Figure 22: t-SNE and PCA visualization of context embedding extracted from trajectories collected in Hopper environment.
@@
Table 5: Sub-standard comparison of average return with baselines in MuJoCo benchmark (over 5 seeds). The blue background indicates that our method surpasses the baseline, and bold text signifies the highest average return. The numerical results for PPO+DOMINO are copied from Mu et al. [2022]; the numerical results for PPO+CaDM, Vanilla+CaDM, and PE-TS+CaDM are copied from Lee et al. [2020].
@@
SaTESAC 911$\pm$54 590$\pm$111 552$\pm$119 7303$\pm$1353 4077$\pm$424 1723$\pm$88
SaCCM 1015$\pm$60 629$\pm$126 588$\pm$92 7662$\pm$316 4384$\pm$223 1828$\pm$52
SlimHumanoid Handcapped Half-Cheetah
@@ Appendix G A comparison with DOMINO and CaDM
In this section, we give a brief comparison between our methods and methods from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] in the MuJoCo benchmark because we are using the exact same environmental setting.
@@
DOMINO [Mu et al., 2022] is based on the InfoNCE $K$-sample estimator. Their implementation, PPO+DOMINO, is a model-free RL algorithm with a pre-trained context encoder. This encoder reduces the demand for large contrastive batch sizes during training by decoupling representation learning for each modality, simplifying tasks while leveraging shared information. However, a pre-trained encoder necessitates a large sample volume, with DOMINO training PPO agents for 5 million timesteps on the MuJoCo benchmark. In contrast, SaTESAC and SaCCM, trained for 1.6 million timesteps without pre-trained encoders, achieve considerably higher average returns across four environments (Table 5). Therefore, it is crucial to focus on extracting mutual information in contrastive learning that directly optimises downstream tasks, integrating rather than segregating representation learning from task performance.
@@
CaDM [Lee et al., 2020] proposes a context-aware dynamics model adaptable to changes in dynamics. Specifically, they utilise contrastive learning to learn context embeddings, and then predict the next state conditioned on them. We copy the numerical results of PPO+CaDM, Vanilla+CaDM, and PE-TS+CaDM from CaDM [Lee et al., 2020] as their environmental setting is identical to ours, where PPO+CaDM is a model-free RL algorithm, while Vanilla+CaDM and PE-TS+CaDM are model-based. The model-free RL approach, PPO+CaDM, is trained for 5 million timesteps on the MuJoCo benchmark. As shown in Table 5, SaTESAC and SaCCM significantly outperform PPO+CaDM. The model-based RL algorithms, Vanilla+CaDM and PE-TS+CaDM, require 2 million timesteps for learning in model-based setups, compared to our fewer samples (i.e., million timesteps). In the Ant environment, Vanilla+CaDM and PE-TS+CaDM achieve higher returns than SaTESAC and SaCCM; similarly, in the Half-cheetah environment, PE-TS+CaDM outperforms them. Results in the SlimHumanoid and Crippled Half-cheetah environments show that skill-aware context embeddings are notably effective. An insight here is that our method outperforms the model-free CaCM approach, but not the model-based one. Therefore, we consider that a model-based approach to SaMI could be an interesting extension for future work.

### v3

@@ D.3 Implementation details
In this section, we provide the implementation details for SaMI. We present the pseudo-code for using SaNCE during meta-training and meta-testing in Algorithms 1 and 2. Our codebase is built on top of the publicly released implementation of Stable Baselines3 by Raffin et al. [2021] and the implementation of InfoNCE by Oord et al. [2019]. A public, open-source implementation of SaMI is available at https://github.com/uoe-agents/SaMI.
Base algorithm. We use SAC [Haarnoja et al., 2018] for the downstream evaluation of the learned context embedding. SAC is an off-policy actor-critic method that leverages the maximum entropy framework for soft policy iteration. At each iteration, SAC performs soft policy evaluation and improvement steps. We use the same SAC implementation across all baselines and other methods. During the meta-training phase, we trained agents for 1.6 million timesteps in each environment on the Panda-gym and MuJoCo benchmarks. For meta-testing, we evaluated 100 episodes in each environment, with tasks randomly sampled from the moderate and extreme task sets.
Table 4: Hyperparameters used in the Panda-gym and MuJoCo benchmarks. Most hyperparameter values remain unchanged across tasks, except for the contrastive batch size and the SaNCE loss coefficient.
@@
Encoder architecture. In our method, the context encoder $\psi$ is modelled as a Long Short-Term Memory (LSTM) network that produces a 128-dimensional hidden state vector, which is subsequently processed through a single-layer feed-forward network to generate a 6-dimensional context embedding. We aim for the agent to complete the three steps of "explore effectively, infer, adapt" within an episode. Therefore, we initialise the hidden state and cell state of the LSTM to zero at the start of each episode. Both the actor and critic use the same context encoder to embed trajectories. For contrastive learning, SaNCE utilises a momentum encoder $\psi^{*}$ to generate positive and negative context embeddings [Laskin et al., 2020, He et al., 2020]. Formally, denoting the parameters of $\psi$ as $\theta_{\psi}$ and those of $\psi^{*}$ as $\theta_{\psi^{*}}$, we update $\theta_{\psi^{*}}$ as follows:
@@
Here $m\in[0,1)$ is a soft-update rate. Only the parameters $\theta_{\psi}$ are updated by back-propagation. The momentum update in Equation (13) makes $\theta_{\psi^{*}}$ evolve more smoothly by having them slowly track the $\theta_{\psi}$ with $m\ll 1$ (e.g., $m=0.05$ in this research). This means that the target values are constrained to change slowly, greatly improving the stability of learning.
@@
Algorithm 1 SaNCE Meta-training
0:  Batch of training tasks $\{e_{n}\}_{n=1,...,N}$ from $\xi_{train}(e)$, soft-update rate $m$;
1:  Initialize RL replay buffer $\mathcal{B}_{RL}$, encoder replay buffer $\mathcal{B}_{enc}$;
2:  Initialize parameters $\psi$ for context encoder, $\psi^{*}$ for momentum context encoder and $\phi$ for the off-policy SAC;
3:  while not done do
4:     for each task $e_{n}$ do
5:        for Roll-out time steps do
6:           for time step $t$ < maximum episode length $T$ do
7:              Update context embedding $c_{n}\sim\psi(c_{n}|\tau_{c_{n},0:t})$
8:              Roll-out policy $\pi_{c_{n}}(a_{t}|s_{t},c_{n})$ and accumulate transition $(s_{t},a_{t},r_{t},s_{t+1})$;
9:           end for
10:           Add trajectory $\tau_{c_{n}}=\{s_{0},a_{0},r_{0},s_{1},r_{1},...,s_{T},a_{T},r_{T}\}$ to replay buffer $\mathcal{B}_{RL}^{n}$ and $\mathcal{B}_{enc}^{n}$;
11:        end for
12:     end for
13:     for each training step do
14:        for each task $e_{n}$ do
15:           Sample RL batch $\{\tau_{c_{n}}\}\sim\mathcal{B}_{RL}^{n}$;
16:           Sample a positive sample $\tau_{c_{n}}^{+}$ for generating query with highest return, positive samples $\{\tau_{c_{n}}^{-}\}$ and negative samples $\{\tau_{c_{n}}^{-}\}$ for encoding positive and negative embeddings;
17:           Update $\phi$ with RL loss $\mathcal{L}_{\text{RL}}$;
18:           Update $\psi$ with SaNCE loss $\mathcal{L}_{\text{SaNCE}}$ and RL loss $\mathcal{L}_{\text{RL}}$;
19:           $\theta_{\psi^{*}}\leftarrow m\cdot\theta_{\psi}+(1-m)\cdot\theta_{\psi^{*}}$;
20:        end for
21:     end for
22:  end while
Algorithm 2 SaNCE Meta-testing
0:  Batch of training tasks $\{e_{n}\}_{n=1,...,N}$ from $\xi_{test}(e)$;
1:  while not done do
2:     for each task $e_{n}$ do
3:        for each episode do
4:           for time step $t$ < maximum episode length $T$ do
5:              Update context embedding $c_{n}\sim\psi(c_{n}|\tau_{c_{n},0:t})$
6:              Roll-out policy $\pi_{c_{n}}(a_{t}|s_{t},c_{n})$ and accumulate transition $(s_{t},a_{t},r_{t},s_{t+1})$;
7:           end for
8:        end for
9:     end for
10:  end while
@@ E.1 Balance contrastive and RL updates: loss coefficient $\alpha$
While past work has optimised hyperparameters to balance the contrastive loss coefficient $\alpha$ relative to the RL objective [Jaderberg et al., 2016, Bachman et al., 2019], we use both the contrastive and RL objectives with equal weight, setting $\alpha=1.0$ for the MuJoCo benchmark and $\alpha=0.01$ for the Panda-gym benchmark. Additionally, we analyse the effect of the loss coefficient $\alpha$ for CCM, SaTESAC, and SaCCM in the MuJoCo (Figure 12) and Panda-gym (Figure 11) benchmarks.
@@ E.2.1 Buffer size
Figure 13: Comparison of different buffer sizes in the MuJoCo benchmark on training tasks (averaged over 5 seeds). Buffer sizes are 400,000, 100,000, 10,000, and 1,000.
@@ E.2.2 Contrastive batch size
Figure 14: Comparison of different contrastive batch sizes in the MuJoCo benchmark on training tasks (averaged over 5 seeds). Contrastive batch sizes are 512, 128, 16, and 8.
@@ F.1.1 Visualisation of context embedding
We visualise the context embedding using UMAP [McInnes et al., 2020] (Figure LABEL:fig:vis_umap_panda) and t-SNE [Van der Maaten and Hinton, 2008] (Figure 16). When the mass of the cube is high (30 kg and 10 kg), the agent learned the Push skill (indicated by the yellow bounding box in Figure 5.1(a)), whereas with lower masses, the agent learned the Pick&Place skill. However, as shown in Figure 16(b), CCM did not display clear skill grouping. This indicates that SaMI extracts high-quality skill-related information from the trajectories and enables agents to autonomously discover a diverse range of skills for handling multiple tasks.
Figure 15: UMAP visualisation of context embeddings extracted from trajectories collected in the Panda-gym environments.
Figure 16: t-SNE visualisation of context embeddings extracted from trajectories collected in the Panda-gym environments.
@@ F.1.2 Heatmap of Panda-gym benchmark
Figure 17: Heatmap of success rates and learned skills of SaCCM. For each grid, the $(i,j)^{\text{th}}$ location shows the probability of the skills executed over 100 evaluations with $(\text{mass}=i,\text{friction}=j)$.
This section presents the heatmap results and further analysis of TESAC, CCM, and SaTESAC on the Panda-gym benchmark. From the heatmap results of SaTESAC and SaCCM (Figure 6), we observe that, with higher cube masses (30 and 10 kg), the agent executed the Push skill (indicated by the clustered points within the yellow bounding box in Figure 5.1(a)). At lower masses, the agent executed the Pick&Place skill.
In contrast, as shown in Figures 17(e-f), CCM predominantly learned the Pick&Place skill, resulting in a drop in success rates for tasks with $mass=30$, as the agent could not lift the cube off the table using the Pick&Place skill, as depicted in Figure 17(d). The visualisation of the context embedding (Figure 16) did not reveal clear skill grouping across different tasks.
Finally, TESAC primarily mastered the Push skill. The Push skill is relatively simpler to learn than the Pick&Place skill, as it does not require the agent to manipulate its fingers to pick up cubes. Consequently, TESAC’s success rate notably decreased in environments with higher friction.
@@ F.2 MuJoCo
SaMI enables RL agents to be versatile and embody multiple skills. Additionally, video demos (available at https://github.com/uoe-agents/SaMI) provide a better demonstration of the different skills.
Figure 18: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in the Crippled Half-Cheetah environment. "cripple_set" refers to the index of the crippled joint. The figure shows the context embeddings of tasks in three moderate test settings. Combined with the video demos2 for skill analysis, the Crippled Half-Cheetah robot executed three distinct forward running skills.
Figure 19: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in the Crippled Ant environment. "cripple_set" refers to the index of the crippled joint. When 3 or 4 legs are available, the Ant robot (trained with SaCCM) rolls to adapt to varying mass and damping. However, with only 2 adjacent legs during zero-shot generalisation, it switches to walking. If 2 opposite legs are available, the Ant can still roll but eventually tips over. Please refer to the video demos2.
Figure 20: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in the Half-Cheetah environment. During zero-shot generalisation, SaCCM demonstrates different skills for running forwards at various speeds, as well as skills for doing flips and faceplanting.
Figure 21: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Ant environment. The Ant robot learned a single skill, rolling, to adapt to different mass and damping values.
Figure 22: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in SlimHumanoid environment. The Humanoid Robot crawls on the ground using one elbow. When the damping is relatively high (damping=1.8), the Humanoid Robot can crawl forward stably, but when the damping is low (damping=0.4), it tends to roll.
Figure 23: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in HumanoidStandup environment. SaCCM and SaTESAC learned a sitting posture that makes it easier to stand up, allowing it to generalise well when mass and damping change.
Figure 24: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Hopper environment. Combined with the video demos 2 for skill analysis, the plots for SaCCM and SaTESAC show two skills: 1) when the mass is low, the Hopper hops in an upright posture; 2) when the mass is higher, the Hopper hops forward on the floor.
Figure 25: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Crippled Hopper environment. "cripple_set" refers to the index of the crippled joint. SaCCM and SaTESAC learned to take a big hop forward at the beginning (i.e., effective exploration) and then switch to different skills based on environmental feedback.
Figure 26: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Walker environment. The Walker learned a single skill, hopping forward on the floor through both right and left legs.
Figure 27: t-SNE and UMAP visualisations of context embeddings extracted from trajectories collected in Crippled Walker environment. "cripple_set" refers to the index of the crippled joint. The Crippled Walker (trained with SaTESAC and SaCCM) learned to hop forward using the right leg (cripple_set:0 and cripple_set:2) in the training and moderate test tasks and switched to hopping forward using the left leg (cripple_set:4) in the extreme test tasks.
Table 5: Environmental features used for MuJoCo benchmark from DOMINO and CaDM.
Training Test (Moderate) Test (Extrame) Episode Length
Crippled Half-cheetah $m\in\{0.75,0.85,1.0,1.15,1.25\}$ $d\in\{0.75,0.85,1.0,1.15,1.25\}$ Crippled Joints: $\{0,1,2,3\}$ $m\in\{0.40,0.50,1.50,1.60\}$ $d\in\{0.40,0.50,1.50,1.60\}$ Crippled Joints: $\{4,5\}$ $m\in\{0.20,0.40,1.60,1.80\}$ $d\in\{0.20,0.40,1.60,1.80\}$ Crippled Joints: $\{4,5\}$ 1000
@@
Table 6: Comparison of average return $\pm$ standard deviation with baselines in MuJoCo benchmark (over 5 seeds). The bold text signifies the highest average return. The numerical results for PPO+DOMINO are copied from Mu et al. [2022]; the numerical results for PPO+CaDM, Vanilla+CaDM, and PE-TS+CaDM are copied from Lee et al. [2020].
@@
SaTESAC 908$\pm$65 640$\pm$117 532$\pm$88 7430$\pm$1026 4058$\pm$890 1780$\pm$102
SaCCM 928$\pm$141 635$\pm$94 555$\pm$88 7154$\pm$965 3849$\pm$689 1926$\pm$218
SlimHumanoid Crippled Half-Cheetah
@@ Appendix G A comparison with DOMINO and CaDM
In this section, we give a brief comparison between our methods and methods from DOMINO [Mu et al., 2022] and CaDM [Lee et al., 2020] in the MuJoCo benchmark because we are using the exact same environmental setting (shown in Table 5).
@@
DOMINO [Mu et al., 2022] is based on the InfoNCE $K$-sample estimator. Their implementation, PPO+DOMINO, is a model-free RL algorithm with a pre-trained context encoder. This encoder reduces the demand for large contrastive batch sizes during training by decoupling representation learning for each modality, simplifying tasks while leveraging shared information. However, a pre-trained encoder necessitates a large sample volume, with DOMINO training PPO agents for 5 million timesteps on the MuJoCo benchmark. In contrast, SaTESAC and SaCCM, trained for 1.6 million timesteps without pre-trained encoders, achieve considerably higher average returns across four environments (Table 6). Therefore, it is crucial to focus on extracting MI in contrastive learning that directly optimises downstream tasks, integrating rather than segregating representation learning from task performance.
@@
CaDM [Lee et al., 2020] proposes a context-aware dynamics model adaptable to changes in dynamics. Specifically, they utilise contrastive learning to learn context embeddings, and then predict the next state conditioned on them. We copy the numerical results of PPO+CaDM, Vanilla+CaDM, and PE-TS+CaDM from CaDM [Lee et al., 2020] as their environmental setting is identical to ours, where PPO+CaDM is a model-free RL algorithm, while Vanilla+CaDM and PE-TS+CaDM are model-based. The model-free RL approach, PPO+CaDM, is trained for 5 million timesteps on the MuJoCo benchmark. As shown in Table 6, SaTESAC and SaCCM significantly outperform PPO+CaDM. The model-based RL algorithms, Vanilla+CaDM and PE-TS+CaDM, require 2 million timesteps for learning in model-based setups, compared to our fewer samples (i.e., million timesteps). In the Ant environment, Vanilla+CaDM and PE-TS+CaDM achieve higher returns than SaTESAC and SaCCM; similarly, in the Half-cheetah environment, PE-TS+CaDM outperforms them. Results in the SlimHumanoid and Crippled Half-cheetah environments show that skill-aware context embeddings are notably effective. An insight here is that our method outperforms the model-free CaCM approach, but not the model-based one. This is consistent with what is empirically found in CaDM [Lee et al., 2020]: prediction models are more effective when the transition function changes across tasks. Therefore, we consider that a model-based approach to SaMI could be an interesting extension for future work.
Appendix H Statistical hypothesis tests (paried t-tests)
We used a t-test [Rice and Rice, 2007] to conduct a statistical hypothesis test to determine whether SaMI brought a statistically significant improvement. we reported the p-value of the t-test in MuJoCo (Table 7) and Panda-gym (Table 8) benchmarks. * next to the number is used to indicate that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI at a significance level of 0.05. From Table 7 and 8, SaMI brings significant improvement on the extreme test set in which the RL agent needs to execute diverse skills. The statistically significant test aligns with our results in the skill analysis (i.e., video demos, t-SNE and UMAP visualisation). In complex environments that require high skill diversity from the RL agent, the statistically significant improvement and higher returns/success rates brought by SaMI are evident.
Table 7: The p-value of the statistical hypothesis tests (paried t-tests) for comparing the effectiveness of SaMI in MuJoCo benchmark (over 5 seeds). * next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI at a significance level of 0.05. The “SaTESAC-TESAC” row indicates the p-value for the return improvement brought by SaMI to TESAC; the “SaCCM-CCM” row indicates the p-value for the return improvement brought by SaMI to CCM.
Crippled Ant Crippled Half-cheetah
SaTESAC-TESAC 0.121 0.109 9.54E-07 0.154 0.0024 0.0889
SaCCM-CCM 0.913 0.108 0.008 0.04 0.307 0.106
Ant Half-cheetah
SaTESAC-TESAC 0.136 0.034 0.021 0.346 0.224 0.004
SaCCM-CCM 0.138 0.275 0.163 0.73 0.791 0.005
SlimHumanoid HumanoidStandup
SaTESAC-TESAC 0.139 0.456 0.006 0.037 0.129 0.003
SaCCM-CCM 0.113 0.059 0.008 0.048 0.027 0.01
Hopper Crippled Hopper
SaTESAC-TESAC 0.747 0.089 0.707 0.459 0.69 0.088
SaCCM-CCM 0.52 0.599 0.969 0.967 0.897 0.0002
Walker Crippled Walker
SaTESAC-TESAC 0.312 0.082 0.003 0.223 0.048 0.028
SaCCM-CCM 0.55 0.541 0.022 0.794 0.079 0.011
Table 8: The p-value of the statistical hypothesis tests (paried t-tests) for comparing the effectiveness of SaMI in Panda-gym benchmark (over 5 seeds). * next to the number means that the algorithm with SaMI has statistically significant improvement over the same algorithm without SaMI at a significance level of 0.05. The “SaTESAC-TESAC” row indicates the p-value for the return improvement brought by SaMI to TESAC; the “SaCCM-CCM” row indicates the p-value for the return improvement brought by SaMI to CCM.
Training Test (moderate) Test (extreme)
SaTESAC-TESAC 0.000260 0.000160 0.001310
SaCCM-CCM 0.000230 0.002190 0.001390